OpenStudy (anonymous):

Consider the functions f(x) = x2 − 13 and g(x) = x + 5. What is the value of (f + g)(3)? −4 4 −32 8

OpenStudy (amorfide):

f(3) + g(3) substitute x=3 into the function f then substitute 3 into the function g then add the answers together

hartnn (hartnn):

what di u get?

OpenStudy (anonymous):

idk i dont understand how to do these

hartnn (hartnn):

u know how to find f(3) ?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

its been a while since i did these and i forgot

OpenStudy (amorfide):

f(x) is a function containing x if you want f(3) you want a function containing 3's so wherever x was originally you would replace them all with 3

hartnn (hartnn):

to get f(3) put x=3 in f(x)

OpenStudy (amorfide):

so for your function f(x) = x2 − 13 you want to replace all x with 3 so you get f(3)=3²-13 then do the same with g then add your answers

OpenStudy (anonymous):

f(2) = (2)2 - 13 and g(2) = 2 + 5?

OpenStudy (anonymous):

i mean three lol

OpenStudy (anonymous):

f(3) = (3)2 - 13 and g(3) = 3 + 5?

OpenStudy (amorfide):

yes that is right

OpenStudy (amorfide):

now get an answer for each function add your answers together

OpenStudy (anonymous):

f(3) = 9 - 13 = f(3) = -4

OpenStudy (anonymous):

g(3) = 3 + 5 = g(3) = 8

OpenStudy (anonymous):

now what do i do???

OpenStudy (amorfide):

f+g -4+8

OpenStudy (anonymous):

wait when i do f+g what do i do with the 3's??

OpenStudy (anonymous):

-4+8=4

OpenStudy (anonymous):

so the answer is 4??